OpenStudy (caitlinnr14):

HELP WILL GIVE MEDAL AND FAN FOR BEST ANSWER Choose the equation below that represents the line passing through the point (-5, 1) with a slope of three halves. y - 5 = three halves(x + 1) y + 1 = three halves(x - 5) y + 5 = -three halves(x - 1) y - 1 = three halves(x + 5)

OpenStudy (caitlinnr14):

thx...but when u explain it it is confusing:( @vinnv226

OpenStudy (anonymous):

Where did I confuse you? You have all the variables, you just need to plug them in. M is the slope, which is 3/2 here. x_1 is the x coordinate of your point, and y_1 is the y coordinate of your point.

OpenStudy (caitlinnr14):

oh ok i got it thx you sry i must hav been misreading it

OpenStudy (caitlinnr14):

so it would be the first one ? @vinnv226

OpenStudy (anonymous):

No, make sure you're plugging in correctly. m=3/2 x_1=-5 y_1=1

OpenStudy (caitlinnr14):

so it would be y - 1 =3/2(x + 5)

OpenStudy (caitlinnr14):

@vinnv226

OpenStudy (anonymous):

Right.

OpenStudy (caitlinnr14):

thank you!

OpenStudy (anonymous):

No problem, glad I could help

OpenStudy (anonymous):

thanks u helped me to